Current StatusOperation Status

2014-2015 Solar Paddle Incident and Pointing Mechanism Switch

Long term status of the TANSO-FTS operation and the L1B product. There is no change in TANSO-CAI operations.

    • May 24, 2014, Rotation of one of the two solar paddles stopped. Still enough power to operate TANSO-FTS (TIR cooler, pointing mechanism), CAI.
    • Jun. 1, 2014, FTS observation restarted. Much Faster ZPD shift >> 800 fringe-biased interferogram. Needs new ILS (ILS800SS) for V161.160 or bias-correction apodization (V200.200). There is frequent spike flag 4 due to improper detection of the center burst.
    • Since Sep. 4, 2014, Along track (AT) pointing settling down has become worse.
    • Sep. 11, 2014, 1 point (Nadir only) + glint + target mode instead of 3 point CT mode. About 30% interferogram has pointing instability.
    • Nov. 17-Dec. 5, 2014, PID control parameter tuning does NOT work (no improvement).
    • Since Dec. 3, 2014, AT stability became worse. Sep. 22-Oct. 5, 2014, AT angle was limited to within ±10 degrees.
    • From Dec. 15, 2014, Backup system test (wrong Level 1 geometry telemetry).
    • Jan. 26, 2015, optics path selector from the primary to the secondary.
    • Feb. 14, 2015, Level 1 data distribution (data with the back-up pointing mechanism).
    • Aug. 30, 2015, Pointing offset and misalignment were both reflected to Target observation requests planning.

2015 Cryocooler Shutdown and Restart

  1. On Aug. 2, 2015, The band 4 detector-temperature increased from 70K to the environment temperature, while the short-wave infrared bands 1, 2 and 3 are operating normally and providing O2, CO2 and CH4 spectra. TIR operations were stopped (switched off the TIR data processor and shut down cooler control electronics).
  2. On Sep. 14, 2015, the cryocooler (pulse tube cooler) restarted the detector cooling. The most probable root cause of the sudden anomaly on Aug. 2, 2015 is a single event.
    The MCT detector temperature reached 70 K and JAXA resumed the ground data processing of all the TANSO-FTS bands (bands 1, 2, 3, and 4) from Sep. 15, 2015 (UT).

ZPD posiotion changing with time: After Cryocooler Restart.

May, 2018 CDMS incident and restart

Long term status of the TANSO-FTS operation and the L1B product.

  1. On 17 May, 2018, GOSAT transferred to the least load mode (LLM) and all the instrument and data recorders were shut down due to the Command and Data Management System (CDMS) incident.The optics were cooled down and cryo cooler shut down.The condition was similar to the one when the rotation of the solar paddle stopped in May 2014.GOSAT restarted observing and communication with the ground stations and the data distribution was also restarted on 29 May 2018.
    After analyzing the trend since the restart, the parameters in the level 1 processing for TANSO-FTS TIR nonlinearity-correction will be updated and the level 1 data will be reprocessed as Version 220 (T.B.D.).
  2. Data missing period;
    TANSO-FTS: 17/05/2018 18:48:25 (UT) - 24/05/2018 22:38:37(UT)
    TANSO-CAI: 17/05/2018 17:12:54 (UT) - 24/05/2018 21:50:20(UT)
    CAM-H8 (FTS field-of-view camera): 17/05/2018 18:48:29 (UT) - 24/05/2018 22:38:41(UT)

Go to TopGo to Top